OpenStudy (anonymous):

How do you find the real number solution of y^3-5y^2=0??

OpenStudy (anonymous):

factoring helps ...

OpenStudy (anonymous):

we know y=0 is a solution, so assume y not equal to 0 y^3 - 5y^2 = 0 y^3 = 5y^2 ; divide by y^2 y = 5 will be another solution.

OpenStudy (anonymous):

there will be one more solution such that: y(y-5)(y-r) = y^3 - 5y^2